Calculate the hydronium ion conc and the ppt of a solution that is 0.500M HCL
-negativity activity correction
-using activity coefficients
1
Expert's answer
2016-06-17T04:54:24-0400
The hydronium ion conc = 0.5M, becouse of HCl is strong acid.
M((H3O+)=17g/mol, m(H3O+)=0.5*17=8.5g
PPM= 8.5g/1000g=0.008500 or 8500ppm

1. The ionic strength of the solution I=0.5*Sum(Ci*Zi2)
I(0.500M HCL)=0.5*(0.5*12+ 0.5*12)=0.5
2. Lgf=-0.5*(Z2)*√I = -0.5*12*√0.5=-0.3535
3. f is activity ratio = 10 -0.3535=0.44
4. Activity or (a)=f*C, hereof a(H3O+)=0.44*0.5=0.22
the hydronium ion conc=0.22

Answer: PPM=8500ppm; the hydronium ion conc=0.22